# The Forest Behind the Tree: Revealing Hidden Smart Home Communication Patterns

### v2

@@ Abstract
The widespread use of Smart Home devices has attracted significant research interest in understanding their behavior within home networks. Unlike general-purpose computers, these devices exhibit relatively simple and predictable network activity patterns. However, previous studies have primarily focused on normal network conditions, overlooking potential hidden patterns that emerge under challenging conditions. Discovering these hidden flows is crucial for assessing device robustness.
This paper addresses this gap by presenting a framework that systematically and automatically reveals these hidden communication patterns. By actively disturbing communication and blocking observed traffic, the framework generates comprehensive profiles structured as behavior trees, uncovering flows that are missed by more shallow methods. This approach was applied to ten real-world devices, identifying 254 unique flows, with over 27% only discovered through this new method. These insights enhance our understanding of device robustness and can be leveraged to improve the accuracy of network security measures.
@@ I Introduction
As IoT devices have very low computing resources, it is a technical challenge to embark them with all the technology necessary for their correct operation. Notoriously, to stay economically competitive, manufacturers tend to expand the user-visible functionalities on the devices’ limited hardware, neglecting transversal aspects such as robustness or security [3, 4]. Ironically, this compromise might impede the initial incentive behind the design of such systems, i.e. user convenience. Indeed, if devices do not provide robustness in the face of network communication instability, the automation system may be unresponsive, penalizing the user. The motivation behind this work is to assess the robustness of Smart Home devices against network instabilities. More precisely, we envision to discover and describe network communication patterns issued by such devices in the case where their default traffic fails.
@@
Due to their goal-oriented nature, Smart Home devices generally exhibit simple, predictable network communication patterns [5]. Consequently, researchers have designed solutions to express their network behavior in a reduced and exhaustive form, a so-called profile. The IETF has even standardized a format for such profiles, the Manufacturer Usage Description (MUD) [6], albeit admittedly relatively limited [7, 8, 9]. Several methods have been presented in the past to automatically create the profile of a device [10, 11, 12]. For this purpose, its network communications are observed over a certain period of time in the wild or under laboratory conditions. Data mining and analysis techniques are then used to extract compact representations of the network traffic and associate them with specific device events. In order to gain a comprehensive view on the communication profile of a device, relatively complex experimental setups are required [13, 14], due to the variety of possible interactions with the device.
@@
We argue that existing approaches to profiling are insufficient for our goal of gaining comprehensive insight into the network behavior of a Smart Home device, as they do not aim at triggering communication patterns that only become visible under certain network conditions, in particular when the default communication mechanism does not succeed. For example, an IoT device might have a list of alternative domain names for its cloud hosted services that it will only contact if the default domain name fails to resolve or the server behind that name does not respond. Ignoring such behaviors leads to a profile, in which some of the device’s communication patterns are missing or incompletely described.
@@
In this paper, we present a framework to comprehensively uncover the communication patterns of a Smart Home device. For each type of interaction with the device, we first observe the traffic flows that occur in an unconstrained network and build a set of flow descriptors from them. Such descriptors contain Internet and transport layer information as well as application layer data, such as domain names. We then iteratively block flows corresponding to certain descriptors and repeat the interaction with the intent of making new flows appear. Deciding which flows to block and the sequence of blocking experiments is not straightforward. Our aim is to discover as many hidden flows as possible but keep the number of experiments under control. To do this efficiently and not end up in an infinite loop where the same patterns appear over and over again, we store the found descriptors in a tree that we traverse using a breadth-first strategy, and prune using an ad hoc heuristic. The result is a set of multi-level, tree-shaped profiles (one for each type of interaction) for the device that describe the device’s default and alternative communication patterns that we then explore to assess the robustness of the devices’ network communications. Our profiles are conceptually similar to the aforementioned ones, e.g. MUD, and can therefore also be used as input configuration for allow-list firewalls. Current state-of-the-art approaches, which do not cover alternative communication paths, might consider the latter as malicious, whereas they are actually part of the device’s intended behavior, only rarer. Profiles obtained from our framework can thus provide more accurate security configurations that do not block the legitimate communications triggered by a robust device.
@@
  • We apply our framework to a testbed network comprising ten off-the-shelf devices, and generate multi-level profiles for 36 unique events, which sum up to a total of 254 unique network flows discovered, of which 70 were “hidden”, i.e., not part of the default communication patterns of the tested devices, accounting for over 27% of the total number of unique flows.
@@
  • We publish the source code of our signature extraction algorithm and our experimental framework. Those will be made available after the anonymous review process.
The rest of the paper is structured as follows. We describe the problem and the scope of our work in Section II. In Section III, our methodology is explained in detail. Its evaluation is presented in Section IV. Section V presents research works related to ours, and positions them compared to ours. Ultimately, the paper concludes in Section VI.
@@ II Problem statement and scope
The solution that we will present in the next sections requires the collection, inspection and filtering of packets sent to and received from the Smart Home device and the device running the companion app. If the traffic is unencrypted, we leverage its payload. We do not analyze or modify the device firmware, the companion app, or the server side software. We also do not try to decrypt network traffic, therefore we do not consider information in encrypted payload. However, numerous research has shown than relying on unencrypted information and packet or flow metadata, can already provide sufficient information to accurately fingerprint IoT devices [12, 11, 15].
@@ III-A General workflow
Fig. 2 gives an overview on the workflow of our methodology to obtain a profile for a Smart Home device under a certain type of interaction. An interaction can require the usage of a triggering source, such as a companion app installed on a smartphone. In practice, a Smart Home device may support different interactions (e.g., for a smart lamp: switching on/off, changing its color, etc.) and triggering sources, and the workflow described here must be repeated for each of them. The workflow consists of multiple steps summarized below.
@@
Step 1 consists in performing the interaction with the device under investigation and recording the network traffic. To this end, we connect the Smart Home device (or, for non-IP devices, their gateway or hub) and the device running the triggering source to a (wireless or wired) LAN and capture the traffic on the Access Point (AP), respectively switch. We call an interaction and the resulting device state an event.
@@ III-B Step 1: Traffic capturing
As our approach is based on blocking traffic, it can happen that the interaction does not cause the desired state change in the Smart Home device. For the event signature extraction described in step 2, it is important to filter out such unsuccessful event executions. We do this by verifying the device state, e.g. whether the lamp has turned on, after each capture. To obtain the device state, we leverage the device’s API. We believe this method is more precise and reliable than the approach followed by Mandalari et al. [23] which compares screenshots of the companion app to check whether the displayed status of the Smart Home device has changed. If there is no API, we fall back to screenshots, too, but we compute the Structural Similarity Measure (SSIM) [24] and consider two screenshots identical if the SSIM is over an empirically defined threshold, instead of directly comparing image pixels like Mandalari et al. The result are $m^{+}\leq m$ traffic captures of successful event executions.
@@ III-C Step 2: Event signature extraction
For each of the $m^{+}$ packet traces, the recorded packets are first aggregated to Flow IDs, and then the event signature is extracted. This happens in multiple steps that we now describe.
@@ III-C1 Replacing IP addresses by domain names
During an event, the Smart Home device or the companion app may communicate with cloud services provided by the device vendor. Where possible, we replace non-local source and destination IP addresses by domain names. This makes the signature extraction more robust against changes in the IP addresses caused by cloud migrations or load balancing. To do so, we keep a table matching encountered domain names with their corresponding IP address(es). We first populate the table with entries from the LAN gateway’s DNS cache, and then update it whenever a packet bearing domain data is observed, i.e. a DNS query/response or a TLS Client Hello message with the SNI extension. [^1] We also tried reverse DNS lookups, but since most of the servers contacted during our experiments were hosted at big cloud providers such as Amazon, no meaningful data was obtained, and we abandoned this approach.
@@ III-D Step 3: Blocking flows with the event signature tree
The event signature tree is a connected, acyclic, rooted tree where the nodes of the tree are Flow IDs, with the exception of the root node. The tree is used to keep track of the Flow IDs already seen and blocked.
@@ III-D1 Tree creation
After obtaining the event signature $s$ from step 2, we add the Flow IDs in the signature as a child node to the root node. For the next iteration of steps 1 through 3, we select a node, i.e. a Flow ID, that has not been visited yet from the tree and instruct the firewall to block all packets matching that node and all its ancestors in the tree. The node will be marked as visited and the newly found Flow IDs will be added as children to it, and the procedure is repeated.
@@ IV-A Experiment setup
We apply our methodology to a controlled testbed network comprised of real-world devices, mimicking a typical Smart Home network, depicted in Fig. 4. We instrument an array of commercial, off-the-shelf Smart Home devices, including four power plugs, three cameras, and three light bulbs. The profiled interactions, per device category, are the following:
@@
For all interactions other than booting, we derive up to three usage scenarios, to account for various ways of controlling the device:
@@
The smartphone running the device-specific apps is a Crosscall CORE-X4 [31], running Android 10.
@@
Kasa Smart toggle 12
TP-Link Tapo toggle 13
SmartThings toggle 17
@@
Tuya Smart toggle 16
@@
SmartThings toggle 29
SmartThings brightness 30
SmartThings color 31
Alecto Smart-Bulb10 (Tuya) [39] / boot 9
Tuya Smart toggle 32
Tuya Smart color 33
Tapo L530E [40] / boot 10
TP-Link Tapo toggle 34
TP-Link Tapo brightness 35
TP-Link Tapo color 36
@@
The complete device references are listed in Table I, along with the corresponding smartphone apps and interactions instrumented. We only consider IP traffic, whether over Ethernet or Wi-Fi. Therefore, for the two devices communicating using Zigbee, we profile the traffic issued by their hub, i.e. the Hue bridge for the Hue lamp, and the SmartThings hub for the SmartThings Outlet, respectively. For the boot interaction experiments over those devices, we synchronously power-cycle both the device itself and its hub.
@@
  • Traffic capture timeout: $d=20$ seconds (90 seconds for event boot);
  • Duration between two event iterations: randomly chosen between 40 (120 for event boot) and 150 seconds.
@@ IV-B Example signature tree
As an example of our framework’s results, we present the event signature tree generated for one device of our corpus, namely the TP-Link HS110 smart plug [26]. This device being a smart power plug, it provides one interaction, i.e. toggling it on and off. For this experiment, we controlled the device using its official Android companion app, i.e. Kasa Smart [41]. The resulting signature tree is displayed in Fig. 5. To avoid unnecessarily cluttering the figure, we only show the first occurrence of each unique flow. In general, every first-level flow occurs as a child of every tree node. In certain cases, a hidden flow will appear at deeper levels, which embodies a backup communication strategy when the default one fails.
@@ IV-C Assessing device event robustness
The experiments in our testbed of devices result in the extraction of 254 unique Flow IDs, of which 70 are hidden ones (i.e. 27.56%). Hidden Flow IDs embody the alternative communication strategies that a device might use if the default one fails, effectively enhancing the device’s robustness to network instability (remember that we only keep Flow IDs associated with successful event executions). As such, we rebrand the count of hidden Flow IDs as the robustness score, and compute it over all the instrumented device events. Out of the 36 instrumented events, 23 (63.9%) have a robustness score of at least one (mean 1.94). This value is encouraging: most types of interactions dispose of at least one backup strategy in the case of a failure. Comprehensive results over all events are shown in the appendix: Fig. 9 shows the count of first-level and hidden Flow IDs; Fig. 10 shows the robustness score.
To gain further insight into the distribution of robustness scores, Fig. 6 shows the same robustness score, grouped along three categories, namely:
@@
We observe that the different device categories do not provide the same robustness. The most robust devices seem to be the power plugs. A rationale might be that, as the plugs’ operation (switching on or off) is the most simple, it takes less effort to the manufacturers to provide backup strategies.
Regarding the companion app, Tuya seems to be less robust than the others. While most apps provide device control either via the local network or through the cloud, Tuya devices are notoriously reluctant to LAN-based control, preferring cloud-based communication. This decision reduces their robustness by design, as, when the cloud endpoints cannot be reached, the device cannot function properly. SmartThings boasts a good score, both as a controlling app and a manufacturer, helped by a very high robustness score for the SmartThings Outlet’s boot interaction. We can also see that the ”other” app category provides no backup strategy whatsoever, suggesting that a user should prefer using trusted and official apps, even if the devices’ API is theoretically open.
@@
Among the manufacturers, the most robust one seems to be Xiaomi, for which the only device in our testbed was a camera. As cameras are usually tied to more critical functions, such as monitoring a home against potential intrusions, Xiaomi followed a respectable rationale here. TP-Link seems quite robust, while Philips, even as a trusted device manufacturer, disappoints in terms of robustness.

### v3

@@ Abstract
The widespread use of Smart Home devices has attracted significant research interest in understanding their behavior within home networks. Unlike general-purpose computers, these devices exhibit relatively simple and predictable network activity patterns. However, previous studies have primarily focused on normal network conditions, overlooking potential hidden patterns that emerge under challenging conditions. Discovering the latter is crucial for assessing device robustness.
This paper addresses this gap by presenting a framework that systematically and automatically reveals these hidden communication patterns. By actively disturbing communication and blocking observed traffic, the framework generates comprehensive profiles structured as behavior trees, uncovering traffic flows that are missed by more shallow methods. This approach was applied to ten real-world devices, identifying 254 unique flows, with over 27% only discovered through this new method. These insights enhance our understanding of device robustness, and the thus obtained profiles provide a more complete description of the network behavior of devices, as needed, for example, for the configuration of security solutions.
@@ I Introduction
As IoT devices have very low computing resources, it is a technical challenge to embark them with all the technology necessary for their correct operation. Notoriously, to stay economically competitive, manufacturers tend to expand the user-visible functionalities on the devices’ limited hardware, neglecting transversal aspects such as robustness or security [3, 4]. Ironically, this compromise might impede the initial incentive behind the design of Smart Home systems, i.e. user convenience. Indeed, if devices do not provide robustness in the face of network communication instability, they may become unresponsive, disrupting the whole system.
The motivation behind this work is to assess the robustness of Smart Home devices against network instabilities. More precisely, we aim to discover and describe network communication patterns issued by such devices in the case where their default traffic fails.
@@
Smart Home devices usually exhibit simple and predictable network communication patterns [5] under ideal conditions. Consequently, researchers have designed solutions to express their network behavior in a compact form, a so-called profile. The IETF has even standardized a format for such profiles, the Manufacturer Usage Description (MUD) [6], albeit admittedly relatively limited [7, 8, 9]. Several methods have been presented in the past to automatically create the profile of a device [10, 11, 12]. For this purpose, its network communications are observed over a certain period of time, in the wild or under laboratory conditions. Typically, relatively complex experimental setups are required [13, 14], due to the variety of possible interactions with the device. Data mining and analysis techniques are then used to extract compact representations of the network traffic and associate them with the device, or even with specific device events.
@@
We argue that existing approaches to profiling are insufficient for our goal of gaining comprehensive insight into the network behavior of Smart Home devices, as they do not aim at triggering communication patterns that only become visible under certain network conditions, in particular when the default communication mechanism does not succeed. For example, an IoT device might have a list of alternative domain names for its cloud hosted services that it will only contact if the default domain name fails to resolve or the server behind that name does not respond; or it can switch from using TCP to UDP (and vice-versa), if one of the protocols is blocked. Ignoring such behaviors leads to profiles, in which some of the device’s communication patterns are missing or incompletely described, making it difficult to grasp if a device is robust with regard to network events.
@@
In this paper, we present a fully automated framework to comprehensively uncover the communication patterns of a Smart Home device. For each type of interaction with the device, it first observes the traffic flows that occur in an unconstrained network and builds a set of flow descriptors from them. Such descriptors contain Internet and transport-layer information, as well as selected, available application-layer data, such as domain names. The system then iteratively blocks flows corresponding to certain descriptors and repeats the interaction with the intent of making new flows appear. The algorithm selecting which flows to block and the sequence of blocking experiments is not straightforward. Our aim is to discover as many hidden flows as possible but keep the number of experiments under control. To do this efficiently and not end up in an infinite loop where the same patterns appear over and over again, we store the found descriptors in a tree that we traverse using a breadth-first strategy, and prune using an ad hoc heuristic. The result is a set of multi-level, tree-shaped profiles (one for each type of interaction) for the device that describe the device’s default and alternative communication patterns that we then explore to assess the robustness of the device’s network communications.
Conceptually, the profiles obtained by our approach can be treated as enhanced versions of the aforementioned ones, such as MUD, as they encompass the base communication patterns, and augment them with the identified hidden patterns. Therefore, they can also be used as input configuration for allow-list firewalls. Current state-of-the-art approaches, which do not cover alternative communication patterns, might consider the latter as malicious, whereas they are actually part of the device’s intended behavior, only rarer. Profiles obtained from our framework can thus provide more accurate security configurations that do not block the legitimate communications of a robust device.
@@
  • We apply our framework to a testbed network comprising ten off-the-shelf devices, and generate multi-level profiles for 36 unique events, which sum up to a total of 254 unique network flows discovered, of which 70 were “hidden”, i.e., not part of the default communication patterns of the tested devices, accounting for over 27% of the total number of unique flows, representing the portion that traditional techniques are likely to miss.
@@
  • We publish the source code of our signature extraction algorithm and our experimental framework, including the captures of the testbed’s traffic. Those will be made available after the anonymous review process.
The rest of the paper is structured as follows. We describe the problem and the scope of our work in Section II. In Section III, our methodology is explained in detail. Its evaluation is presented in Section IV. Section V presents related work and positions them compared to ours. Ultimately, the paper concludes in Section VI.
@@ II Problem statement and scope
The solution that we will present in the next sections requires the collection, inspection and filtering of packets sent to and received from the Smart Home device and the device running the companion app. If the traffic is unencrypted, we leverage its payload. We do not analyze or modify the device firmware, the companion app, or the server side software. We also do not try to decrypt network traffic, therefore we do not consider information in encrypted payload. However, numerous research has shown that relying on unencrypted information and packet or flow metadata, can already provide sufficient information to accurately fingerprint IoT devices [12, 11, 15].
The terms IoT and Smart Home are quite fuzzily defined in the literature. Our work focuses on domestic devices, augmented with IoT capabilities, which exhibit goal-oriented, easily identifiable functions, e.g. power plugs, light bulbs, and cameras. Devices such as Smart TVs and speaker hubs are considered out of scope, as their capabilities have broadened tremendously over the years, including the possibility to install third-party apps to add new features. This makes the network behavior of such devices very complex and customizable, rendering an automated analysis with our framework practically unfeasible.
@@ III-A General workflow
Fig. 2 gives an overview on the workflow of our methodology to obtain a profile for a Smart Home device under a certain type of interaction. An interaction can require the usage of a triggering source, such as a companion app installed on a smartphone. In practice, a Smart Home device may support different interactions (e.g., for a smart lamp: switching on/off, changing its color, etc.) and triggering sources, and the workflow described here must be repeated for each of them. The workflow consists of multiple automated steps summarized below.
@@
Step 1 consists in performing the interaction with the device under investigation and recording the network traffic. To this end, we connect the Smart Home device (or, for non-IP devices, their gateway or hub) and the device running the triggering source to a wireless (resp. wired) LAN and capture the traffic on the Access Point (AP) (resp. switch). We call an interaction and the resulting device state an event.
@@ III-B Step 1: Traffic capturing
As our approach is based on blocking traffic, it can happen that the interaction does not cause the desired state change in the Smart Home device. For the event signature extraction described in step 2, it is important to filter out such unsuccessful event executions. We do this by verifying the device state, e.g. whether the lamp has turned on, after each capture. To obtain the device state, we leverage third-party libraries designed to control the device, e.g. python-kasa [23] for the TP-Link HS110 plug, while ensuring this does not interfere with the traffic related to the studied interaction. We believe this method is more precise and reliable than the approach followed by Mandalari et al. [24] which compares screenshots of the companion app to check whether the displayed status of the Smart Home device has changed. If we did not find such a library, or we failed to include it in our framework, we fall back to screenshots, too, but we compute the Structural Similarity Measure (SSIM) [25] and consider two screenshots identical if the SSIM is over an empirically defined threshold, instead of directly comparing image pixels like Mandalari et al. The result are $m^{+}\leq m$ traffic captures of successful event executions.
@@ III-C Step 2: Event signature extraction
For each of the $m^{+}$ packet traces obtained in step 1, the recorded packets are first aggregated to Flow IDs, and then the event signature is extracted. This happens in multiple steps that we now describe.
@@ III-C1 Replacing IP addresses by domain names
During an event, the Smart Home device or the companion app may communicate with cloud services provided by the device vendor or other external hosts. Where possible, we replace non-local source and destination IP addresses by domain names. This makes the signature extraction more robust against changes in the IP addresses caused by cloud migrations or load balancing. To do so, we keep a table matching encountered domain names with their corresponding IP address(es). We first populate the table with entries from the LAN gateway’s DNS cache, and then update it whenever a packet bearing domain data is observed, i.e. a DNS query/response or a TLS Client Hello message with the SNI extension. [^1] We also tried reverse DNS lookups, but since most of the servers contacted during our experiments were hosted at big cloud providers such as Amazon, no meaningful data was obtained, and we abandoned this approach.
@@ III-D Step 3: Blocking flows with the event signature tree
The event signature tree is a connected, acyclic, rooted tree where the nodes of the tree are Flow IDs, except for the root node. The tree is used to keep track of the Flow IDs already seen and blocked.
@@ III-D1 Tree creation
After obtaining the event signature $s$ from step 2, we add each Flow ID in the signature as a child node to the root node. For the next iteration of steps 1 through 3, we select a node, i.e. a Flow ID, that has not been visited yet from the tree and instruct the firewall to block all packets matching that node and all its ancestors in the tree. The node will be marked as visited and the newly found Flow IDs will be added as children to it, and the procedure is repeated.
@@ IV-A Experiment setup
We apply our methodology to a controlled testbed network comprised of real-world devices, mimicking a typical Smart Home network, depicted in Fig. 4. We examine an array of commercial, off-the-shelf Smart Home devices, including four power plugs, three cameras, and three light bulbs. The profiled interactions, per device category, are the following:
@@
In a real Smart Home network, a user could choose among various apps to control their devices, including the device’s official companion app (from its manufacturer or vendor), a third-party app, or a Smart Home automation platform which provides unified control over devices from different manufacturers. To cover such cases, for all interactions other than booting, we derive up to three usage scenarios:
@@
The smartphone running the device-specific apps is a Crosscall CORE-X4 [32] with the Android 10 OS.
@@
Kasa Smart toggle 11
TP-Link Tapo toggle 12
SmartThings toggle 13
@@
Tuya Smart toggle 17
@@
Alecto Smart-Bulb10 (Tuya) [40] / boot 9
Tuya Smart toggle 29
Tuya Smart color 30
Tapo L530E [41] / boot 10
TP-Link Tapo toggle 31
TP-Link Tapo brightness 32
TP-Link Tapo color 33
SmartThings toggle 34
SmartThings brightness 35
SmartThings color 36
@@
The complete device references are listed in Table I, along with the corresponding smartphone apps and interactions instrumented. As explained, we only consider IP traffic, whether over Ethernet or Wi-Fi. Therefore, for the two devices communicating using Zigbee, we profile the traffic issued by their hub, i.e. the Hue bridge for the Hue lamp, and the SmartThings hub for the SmartThings Outlet, respectively. For the boot interaction experiments over those devices, we synchronously power-cycle both the device itself and its hub.
@@
  • Traffic capture timeout: $d=20$ seconds (90 seconds for interaction boot);
  • Duration between two event iterations: randomly chosen between 40 (120 for interaction boot) and 150 seconds.
@@ IV-B Example signature tree
As an example of our framework’s results, we present the event signature tree generated for one device of our corpus, namely the TP-Link HS110 [27]. This device being a smart power plug, it provides one interaction, i.e. toggling it on and off. For this experiment, we controlled the device using its official Android companion app, i.e. Kasa Smart [42]. The resulting signature tree is displayed in Fig. 5. To avoid unnecessarily cluttering the figure, we only show the first occurrence of each unique flow. We empirically observe that each one of the listed first-level flows also occurs as a child of every other node in the tree, i.e. the communication pattern is relatively stable. In certain cases, a hidden flow will appear at deeper levels, which embodies a backup communication strategy when the default one fails.
@@ IV-C Assessing device event robustness
The experiments in our testbed of devices result in the extraction of 254 unique Flow IDs, of which 70 are hidden (i.e. 27.56%). Their distribution over the 36 tested event configurations is shown in Fig. 11 in the appendix. Hidden Flow IDs embody the alternative communication strategies that a device might use if the default one fails, effectively enhancing the device’s robustness to network instability (remember that we only keep Flow IDs associated with successful event executions). As such, we rebrand the count of hidden Flow IDs as the robustness score, and compute it for each tested event; Fig. 6 shows the results.
Figure 6: Robustness score for all events.
Out of the 36 instrumented events, 23 (63.9%) have a robustness score of at least one (mean 1.94). This value is encouraging, as it means that most types of interactions dispose of at least one backup strategy in the case of a failure.
To gain further insight into the distribution of robustness scores, Fig. 7 shows the same robustness score, grouped along three axes, namely:
@@
We identify an event with a robustness score significantly higher than all others: the boot event for the SmartThings Outlet device. Being a Zigbee-communicating device, its boot event consists in switching on the plug itself as well as its related hub. As hubs are commonly used to provide centralized control over a Smart Home network, they must provide a way to detect the devices present in the network. It is expected that, if the default way fails, the hub will follow backup communication paths to provide this functionality.
We observe that the different device categories in our testbed do not provide the same robustness. The most robust devices seem to be the power plugs. A rationale might be that, as the plugs’ operation (switching on or off) is the simplest, it takes less effort to the manufacturers to provide backup strategies. However, given the size of the testbed compared to the vast number of devices on the market, this observation should be taken as indicative only. This also applies to the following discussion of the companion apps and manufacturers.
Among the companion apps, Tuya seems to be less robust than the others. While most apps provide device control either via the local network or through the cloud, Tuya devices are notoriously reluctant to LAN-based control, preferring cloud-based communication. This decision reduces their robustness by design, as, when the cloud endpoints cannot be reached, the device cannot function properly. SmartThings boasts a good score, both as a controlling app and a manufacturer, helped by a very high robustness score for the SmartThings Outlet’s boot interaction. We can also see that the ”other” app category provides no backup strategy whatsoever, suggesting that a user should prefer using trusted and official apps, even if the devices’ API is theoretically open.
@@
The Xiaomi camera also stands out with a high mean score. As cameras are usually tied to more critical functions, such as monitoring a home against potential intrusions, Xiaomi followed a respectable rationale here. TP-Link seems quite robust, while Philips, even as a trusted device manufacturer, disappoints in terms of robustness.